OpenStudy (anikate):

Decide whether the sequence is an arithmetic, Geometric, or neither. Explain your answer.

OpenStudy (anikate):

question 1.... 6,24,96,384

OpenStudy (anonymous):

I think that question is related to another question.

OpenStudy (anikate):

the post was the instruction

OpenStudy (anonymous):

it is a multiple of 4 from 6 on that is simple arithmetic

OpenStudy (anonymous):

6*4 = 24 24*4 = 96 96*4 = 384 etc...